# CAMPOS II. The onset of protostellar disk substructures and planet formation

### v1

@@
The 1.3 mm CAMPOS survey has resolved 90 protostellar disks with $\sim$15 au resolution across the Ophiuchus, Corona Australis, and Chamaeleon star-forming regions. To address the fundamental question, ‘When does planet formation begin?’, we combined the CAMPOS sample with literature observations of Class 0-II disks (bolometric temperature, $T_{\rm bol}\leq 1900$ K), all mapped at 1.3 mm with resolutions ranging from 4 to 33 au. To investigate substructure detection rates as a function of bolometric temperature, we restricted the sample to disks observed at the 1.3 mm wavelength, with inclinations below 75, linear resolution $\leq 20$ au and resolved with at least four resolution elements ($\theta_{\rm disk}/\theta_{\rm res}\geq 4$). We also considered the effects of extinction correction and the inclusion of Herschel Space Telescope data on the bolometric temperature measurements, to constrain the lower and upper limits of bolometric temperature for each source. We find that by $T_{\rm bol}\sim 200-400\,K$, substructure detection rates increased sharply to $\sim$60 %, corresponding to an approximate age of $0.2-0.4$ Myr. No substructures are detected in Class 0 disks. The ratio of disk-averaged brightness temperature to predicted dust temperature shows a trend of increasing values toward the youngest Class 0 disks, suggesting higher optical depths in these early stages. Classifying disks with substructures into those with and without large central cavities, we find both populations coexisting across evolutionary stages, suggesting they are not necessarily evolutionarily linked. Our statistical analysis confirms that substructures similar to those in Class II disks are already common by the Class I stage, and the emergence of these structures at $T_{\rm bol}\sim 200-400\,K$ could represent only an upper limit.
@@ 2.1 Observational data and the sample construction
Besides the effective angular resolution and linear resolution, it is also important to compare observations at the same wavelength. The thermal and scattered continuum emission at a wavelength $\lambda$ is dominated by grains with a size of $\sim\lambda/2\pi$ (Mie, 1908), meaning ALMA 1.3 mm observations, such as the CAMPOS survey as well as DSHARP and eDISK, are more efficient at probing few 100 $\mu$m sized grains. In contrast, disk observations in the optical/near-infrared mainly probe much smaller sub-micrometer particles. To ensure a consistent survey probing the substructures from similar solids, we only include protostellar disks observed at 1.3 mm wavelength by ALMA. Finally, the occurrence rate of substructures is also influenced by the disk’s inclination angle. Due to the increasing optical depth and reduced emitting size, searching for disk substructures in edge-on disks is extremely challenging. While most of the high-resolution observations towards the more evolved Class II disks avoided highly inclined sources (e.g., DSHARP), the population of Class 0/I disks in nearby regions is more limited, and thus samples of very high-resolution observations (e.g., eDisk) contain a significant amount of highly inclined disk. To take this into account, we exclude all disks in the sample with inclination angles larger than 75.222We assume thin circular disks and use the major-to-minor axis ratio of the dust continuum data to compute the inclination angle ($i$), that is $i=sin^{-1}(R_{\rm min}/R_{\rm maj})\times 180^{\circ}/\pi$.
@@ 2.2 Substructure criteria
We note that following the above criteria, all of the disks considered as with substructures in this work have been already confirmed as such in separate published papers (CAMPOS: Hsieh et al. (2024), eDisk: Ohashi et al. (2023), ODESIA: Cieza et al. (2019, 2021), DSHARP: Andrews et al. (2018a).), and all correspond to substructures that are immediately visible in the 1.3 mm image. Sources that show small variations as a function of distance in $uv-$space or small variation in the intensity radial profile but appear smooth in the image plane counted here as disk without substructures (e.g., Michel et al. 2023). This is a conservative approach, and we may miss some sources with borderline detections, introducing a bias against younger disks with smaller and lower-contrast (shallower) substructures (e.g., Maureira et al. 2024a). The implications of this will be further discussed in Section 5.
@@ 3.3 Bolometric temperature as a disk evolutionary tracer
We cross-matched all protostellar disks with the Young Stellar Objects catalog from the Spitzer Space Telescope “cores to disks” (c2d) and “Gould Belt” (GB) Legacy surveys (Dunham et al., 2015), as well as the Extension of HOPS Out to 500 ParSecs (eHOPS) catalog (Riwaj Pokhrel, private communication), to obtain the bolometric temperature ($T_{\rm bol}$), which serves as a proxy for relative evolutionary age for embedded protostellar systems. The eHOPS catalog contains 1-850 $\mu$m Spectral Energy Distributions (SEDs) assembled from Two Micron All Sky Survey (2MASS), Spitzer, Herschel Space Telescope (Herschel), Wide-field Infrared Survey Explorer (WISE), and James Clerk Maxwell Telescope (JCMT)/SCUBA-2 data. The first eHOPS paper for the Serpens and Aquila molecular clouds was published by Pokhrel et al. (2023). For all other clouds, the SED and protostellar system properties are available in the NASA/IPAC Infrared Science Archive. For sources without a counterpart in both eHOPS and Dunham et al. (2015), we adopted the $T_{\rm bol}$ from the literature. Given that we want to investigate the emergence of substructures, we only included samples with $T_{\rm bol}\leq 1900$ K. The final list of sources with $T_{\rm bol}\leq 1900$ K and an inclination angle less than $75^{\circ}$ is shown in Table LABEL:table:substructure_data. In Table LABEL:table:CAMPOS_Tbol, we present the alternative names for each source in the CAMPOS survey for the Corona Australis, Chamaeleon I & II, Ophiuchus North, and Ophiuchus sources. The cross-matched results for the ODISEA survey are presented in Table LABEL:table:ODISEA2019_Tbol.
@@
In this paper, we consider the bolometric temperature measurements from both the eHOPS catalog and the Dunham et al. (2015) catalog to trace the onset of protostellar disk evolution. Compared to Dunham et al. (2015), the eHOPS catalog provides more complete spectral energy distributions (SEDs) from $1$ to $850\,\mu$m. However, the infrared portion of the SED is not corrected for extinction (Pokhrel et al., 2023). To maintain consistency with the eHOPS catalog, we use the bolometric temperatures without extinction correction from Dunham et al. (2015) for sources without a counterpart in eHOPS. Extinction correction increases the infrared flux and increases the bolometric temperature. Therefore, results based on the eHOPS bolometric temperatures represent a lower age limit for the onset of disk substructures. On the other hand, to investigate the upper age limit for the onset of disk substructures, we adopt the extinction-corrected bolometric temperature from Dunham et al. (2015). The bolometric temperatures were derived from the SED fitting without Herschel Space Observatory data ($70-500\,\mu$m) (Dunham et al., 2015). The extinction correction applied to the infrared data increases the infrared flux and thus further increases the bolometric temperature (Dunham et al., 2015). Results based on the extinction-corrected bolometric temperature from Dunham et al. (2015) represent the upper age limit for the onset of disk substructures as longer-wavelength Herschel data are not included from the SED fitting to compute $T_{\rm bol}$.
@@
The bolometric temperature measured from the eHOPS catalog and Dunham et al. (2015) differs from source to source. The decrease in $T_{\rm bol}$ is not uniformly large when including Hershel data into the Dunham et al. (2015). It depends both on the shape of the SED and on how well the SED was sampled by Dunham et al. (2015). While Herschel photometry was not available for the SED fits in Dunham et al. (2015), many of the protostars had 450 $\mu$m observations from SCUBA, some had 350 $\mu$m observations from SHARC-II on Caltech Submillimeter Observatory, and some brighter objects had Infrared Astronomical Satellite (IRAS) or Infrared Space Observatory (ISO) far-infrared photometry. Spitzer 70 $\mu$m photometry was also available for nearly the full Spitzer-GB sample, but at a factor of 3 lower spatial resolution than Herschel, resulting in confusion for some closer protostars. As a result, some $T_{\rm bol}$ values change by a large amount, while some change by a very small amount. We present the $T_{\rm bol}$ from both eHOPS and Dunham et al. (2015) to quantify the range for the onset of protostellar disk substructures.
@@ 4.1 Protostellar disk substructures versus $T_{\rm bol}$ and disk properties
Figure 1: Protostellar disk substructure detections as a function of disk radius, mass, and bolometric temperature. The latter is taken from the eHOPS catalog. The gray dashed lines indicate the boundary between the Class 0 and Class I, and between the Class I and Class II evolutionary phases (Chen et al., 1995), along with the corresponding expected duration of each Class (Dunham et al., 2015). While the $T_{\rm bol}$ is not linear with age, in general the Class 0 sources have $T_{\rm bol}<70$ K, Class I sources have $65070<T_{\rm bol}<$ 650 K, and Class II and Class III sources have $T_{\rm bol}>650$ K (Bouvier et al., 2021). The black arrow is used to indicate that the particular source has also been classified as Class II in the literature (Narayanan et al., 2023).
@@
Figure 2: Substructure detection as a function of effective resolution, 1.3 mm flux and bolometric temperature, the latter from the the eHOPS catalog. The black arrow is used to indicate that the particular source has also been classified as Class II in the literature (Narayanan et al., 2023).
@@
Figure 3: The upper limit of the onset of disk substructures. Same as Figure 1, but the bolometric temperature is derived from the SED fitting without the Herschel Space Observatory data ($70-500\,\mu$m) (Dunham et al., 2015). Extinction correction is applied to the infared data, which further increases the bolometric temperature (Dunham et al., 2015). The effective resolution and the ALMA Band 6 1.3 mm flux of protostellar disks are color-coded and shown in different symbols according to the presence or absence of substructure. Compact disks with diameters smaller than 4 beams are indicated by crosses. The black arrow is used to indicate that the particular source has also been classified as Class II in the literature (Narayanan et al., 2023).
@@
Figure 4: The upper limit of the onset of disk substructures. Same as Figure 2, but the bolometric temperature is derived from the SED fitting without the Herschel Space Observatory data ($70-500\,\mu$m) (Dunham et al., 2015). Extinction correction is applied to the infared data, which further increases the bolometric temperature (Dunham et al., 2015). The crosses represent compact disks that are not well-resolved with diameters less than 4 times the angular resolution of the observation (beams). The lower limit of the onset of the disk substructures at 200 K found in Figure 1 is also labeled. The black arrow is used to indicate that the particular source has also been classified as Class II in the literature (Narayanan et al., 2023).
@@
Figure 1 illustrates the radius and mass of the protostellar disks as a function of bolometric temperature ($T_{\rm bol}$) derived from the eHOPS catalog. The black dashed lines indicate the $T_{\rm bol}$ that marks the boundary between the Class 0 and Class I, and between the Class I and Class II evolutionary phases (Chen et al., 1995), along with the corresponding expected duration of each Class (Dunham et al., 2015). For reference, we also include the sources that do not comply with the effective resolution criteria ($\theta_{D}/\theta_{\rm res}\geq 4$). The latter sources have, in general, smaller sizes and fluxes/masses.
@@
We find no detection of substructure for all the Class 0 disks in the sample. Figure 2 shows the effective resolution and flux of the disks with and without substructures as a function of bolometric temperature, color-coded by the linear resolution. We can observe that the lack of substructure in Class 0 disks is true across resolutions from 4 to 18 au.
Figure 3 and 4 show similar plots as Figure 1 and 2 but considering the upper limit to bolometric temperatures from Dunham et al. (2015). In that case, the earliest substructures discussed above correspond to $T_{\rm bol}$ of 163 and 300 K, for IRAS 04169+2702555For IRAS 04169+2702, the $T_{\rm bol}$ is adopted from Ohashi et al. (2023). and CFHTWIR-Oph79, respectively.
@@ 4.2 Disk substructure detection rates
Figure 5: Disk substructure detection rates versus the evolutionary stage traced by bolometric temperature. The blue and orange represent the lower and upper limits with the bolometric temperature ($T_{\rm bol}$) measured by the eHOPS (with Herschel, but not extinction corrected) and by Dunham et al. (2015) (no Herschel, but extinction corrected), respectively. The bin size of 8 disks is used to compute the substructure detection rates. The shaded region represents the $1\sigma$ uncertainty from counting statistics. All protostellar disks are observed by ALMA Band 6 at 1.3 mm wavelength, have a disk diameter greater than 4 times the resolution ($\theta_{D}/\theta_{res}\geq 4$), with an inclination angle less than 75 , and linear resolution less than 20 au.
We present the protostellar disk substructure detection rates as a function of bolometric temperature in Figure 5. Each disk is assigned a value of 1 if it exhibits substructures and 0 if it does not. We compute a moving average with a window size of 8, where the kernel slides through $T_{\rm bol}$ and calculates the mean for every group of 8 disks. This mean represents the substructure detection rate. Due to the fixed window size, the detection rates are sampled at discrete levels of 0/8, 1/8, 2/8, …, up to 8/8 — effectively producing 8 bins in the substructure detection rate space.
All protostellar disks in Figure 5 are observed by ALMA Band 6 at 1.3 mm wavelength, have a disk diameter greater than 4 times the resolution ($\theta_{D}/\theta_{res}\geq 4$), with an inclination angle less than 75 , and linear resolution less than 20 au. We adopt the moving average with a window size of 8 to ensure enough sampling resolution across $T_{\rm bol}$ and substructure detection rate space. The blue and orange represent the case when the bolometric temperature is adopted from eHOPS catalog or Dunham et al. (2015). The shaded region represents the uncertainty from the binomial counting uncertainty ($\sigma_{b}$), which is calculated as
@@
Figure 6: From our CAMPOS survey, we divide the disk substructures based on the presence or absence of a central cavity (Hsieh et al., 2024). The “Cavity-type” disk substructure, or “C-type” shows a large central cavity. The second type of disk substructures is the “Centrally-filled type” disk substructure, or “F-type”. Both types of disk substructures can be detected throughout the Class I to Class II phases.
Figure 7: Most of the multiple-ring systems are centrally filled (Type F) (Andrews et al., 2018a; Cieza et al., 2021; Ohashi et al., 2023; Hsieh et al., 2024; Cieza et al., 2019; ALMA Partnership et al., 2015; Stephens et al., 2023). The lack of a large central cavity ($\geq$ 25 % of the disk diameter) implies that a centrally filled disk is a general feature of multiple-ring systems. The cyan circle represents the beam size. The white lines mark the scale of 50 au. Note that HD 163296, RU Lup, AS 209, GW Lup, HD 142666, HD 143006, and Sz 129 either do not have reliable bolometric temperature ($T_{\rm bol}$) measurements or their $T_{\rm bol}\geq 1900$ K, thus are not included in this study.

### v4

@@
The 1.3 mm CAMPOS survey has resolved 90 protostellar disks with $\sim$15 au resolution across the Ophiuchus, Corona Australis, and Chamaeleon star-forming regions. To address the fundamental question, ‘When does planet formation begin?’, we combined the CAMPOS sample with literature observations of Class 0-II disks (bolometric temperature, $T_{\rm bol}\leq 1900$ K), all mapped at 1.3 mm with resolutions ranging from 4 to 33 au. To investigate substructure detection rates as a function of bolometric temperature, we restricted the sample to disks observed at the 1.3 mm wavelength, with inclinations below 75, linear resolution $\leq 20$ au and resolved with at least four resolution elements ($\theta_{\rm disk}/\theta_{\rm res}\geq 4$). We also considered the effects of extinction correction and the inclusion of Herschel Space Telescope data on the bolometric temperature measurements to constrain the lower and upper limits of bolometric temperature for each source. We find that by $T_{\rm bol}\sim 200-400\,K$, substructure detection rates increased sharply to $\sim$60 %, corresponding to an approximate age of $0.2-0.4$ Myr. No substructures are detected in Class 0 disks. The ratio of disk-averaged brightness temperature to predicted dust temperature shows a trend of increasing values toward the youngest Class 0 disks, suggesting higher optical depths in these early stages. Our statistical analysis confirms that substructures similar to those in Class II disks are already common by the Class I stage, and the emergence of these structures at $T_{\rm bol}\sim 200-400\,K$ could represent only an upper limit. Classifying disks with substructures into those with and without large central cavities, we find both populations coexisting across evolutionary stages, suggesting they are not necessarily evolutionarily linked. If protostellar disk substructures do follow an evolutionary sequence, then our results imply that disk substructures evolve very rapidly and thus can be present in all Class I/II stages and/or that they can be triggered at different times.
@@ 2.1 Observational data and the sample construction
Besides the effective angular resolution and linear resolution, it is also important to compare observations at the same wavelength. The thermal and scattered continuum emission at a wavelength $\lambda$ is dominated by grains with a size of $\sim\lambda/2\pi$ (Mie, 1908), meaning ALMA 1.3 mm observations, such as the CAMPOS survey as well as DSHARP and eDISK, are more efficient at probing few 100 $\mu$m sized grains. In contrast, disk observations in the optical/near-infrared mainly probe much smaller sub-micrometer particles. To ensure a consistent survey probing the substructures from similar solids, we only include protostellar disks observed at 1.3 mm wavelength by ALMA. Finally, the occurrence rate of substructures is also influenced by the disk’s inclination angle. Due to the increasing optical depth and reduced emitting size, searching for disk substructures in edge-on disks is extremely challenging. While most of the high-resolution observations towards the more evolved Class II disks avoided highly inclined sources (e.g., DSHARP), the population of Class 0/I disks in nearby regions is more limited, and thus samples of very high-resolution observations (e.g., eDisk) contain a significant amount of highly inclined disks. To take this into account, we exclude all disks in the sample with inclination angles larger than 75.222We assume thin circular disks and use the major-to-minor axis ratio of the dust continuum data to compute the inclination angle ($i$), that is $i=sin^{-1}(R_{\rm min}/R_{\rm maj})\times 180^{\circ}/\pi$.
@@ 2.2 Substructure criteria
We note that following the above criteria, all of the disks considered as with substructures in this work have already been confirmed as such in separate published papers (CAMPOS: Hsieh et al. (2024), eDisk: Ohashi et al. (2023), ODISEA: Cieza et al. (2019, 2021), DSHARP: Andrews et al. (2018a).), and all correspond to substructures that are immediately visible in the 1.3 mm image. Sources that show small variations as a function of distance in $uv-$space or small variation in the intensity radial profile but appear smooth in the image plane are counted here as disks without substructures (e.g., Michel et al. 2023). This is a conservative approach, and we may miss some sources with borderline detections, introducing a bias against younger disks with smaller and lower-contrast (shallower) substructures (e.g., Maureira et al. 2024a). The implications of this will be further discussed in Section 5.
@@ 3.3 Bolometric temperature as a disk evolutionary tracer
We cross-matched all protostellar disks with the Young Stellar Objects catalog from the Spitzer Space Telescope “cores to disks” (c2d) and “Gould Belt” (GB) Legacy surveys (Dunham et al., 2015), as well as the Extension of HOPS Out to 500 ParSecs (eHOPS) catalog (Riwaj Pokhrel, private communication), to obtain the bolometric temperature ($T_{\rm bol}$), which serves as a proxy for relative evolutionary age for embedded protostellar systems (see Appendix B). The eHOPS catalog contains 1-850 $\mu$m Spectral Energy Distributions (SEDs) assembled from Two Micron All Sky Survey (2MASS), Spitzer, Herschel Space Telescope (Herschel), Wide-field Infrared Survey Explorer (WISE), and James Clerk Maxwell Telescope (JCMT)/SCUBA-2 data. The first eHOPS paper for the Serpens and Aquila molecular clouds was published by Pokhrel et al. (2023). For all other clouds, the SED and protostellar system properties are available in the NASA/IPAC Infrared Science Archive444 https://irsa.ipac.caltech.edu/data/Herschel/eHOPS/overview.html. For sources without a counterpart in both eHOPS and Dunham et al. (2015), we adopted the $T_{\rm bol}$ from the literature. Given that we want to investigate the emergence of substructures, we only included samples with $T_{\rm bol}\leq 1900$ K. The final list of sources with $T_{\rm bol}\leq 1900$ K and an inclination angle less than $75^{\circ}$ is shown in Table LABEL:table:substructure_data. In Table LABEL:table:CAMPOS_Tbol, we present the alternative names for each source in the CAMPOS survey for the Corona Australis, Chamaeleon I & II, Ophiuchus North, and Ophiuchus sources. The cross-matched results for the ODISEA survey are presented in Table LABEL:table:ODISEA2019_Tbol.
@@
In this paper, we consider the bolometric temperature measurements from both the eHOPS catalog and the Dunham et al. (2015) catalog to trace the onset of protostellar disk evolution. Compared to Dunham et al. (2015), the eHOPS catalog provides more complete spectral energy distributions (SEDs) from $1$ to $850\,\mu$m. However, the infrared portion of the SED is not corrected for extinction (Pokhrel et al., 2023). To maintain consistency with the eHOPS catalog, we use the bolometric temperatures without extinction correction from Dunham et al. (2015) for sources without a counterpart in eHOPS. The extinction correction applied to the infrared data increases the infrared flux and thus further increases the bolometric temperature (Dunham et al., 2015). Results based on this correction represent the upper age limit for the onset of disk substructures as longer-wavelength Herschel data are not included from the SED fitting to compute $T_{\rm bol}$. Therefore, results based on the eHOPS bolometric temperatures represent a lower age limit for the onset of disk substructures. On the other hand, to investigate the upper age limit for the onset of disk substructures, we adopt the extinction-corrected bolometric temperature from Dunham et al. (2015).
@@
The bolometric temperature measured from the eHOPS catalog and Dunham et al. (2015) differs from source to source. The decrease in $T_{\rm bol}$ is not uniform when including Hershel data into the Dunham et al. (2015). It depends both on the shape of the SED and on how well the SED was sampled by Dunham et al. (2015). While Herschel photometry was not available for the SED fits in Dunham et al. (2015), many of the protostars had 450 $\mu$m observations from SCUBA, some had 350 $\mu$m observations from SHARC-II on Caltech Submillimeter Observatory, and some brighter objects had Infrared Astronomical Satellite (IRAS) or Infrared Space Observatory (ISO) far-infrared photometry. Spitzer 70 $\mu$m photometry was also available for nearly the full Spitzer-GB sample, but at a factor of 3 lower spatial resolution than Herschel, resulting in confusion for protostars with close neighbors. As a result, some $T_{\rm bol}$ values change by a large amount ($\sim$40 %), while some change by a very small amount ($\lesssim$10 %). We present the $T_{\rm bol}$ from both eHOPS and Dunham et al. (2015) to quantify the range for the onset of protostellar disk substructures in Table LABEL:table:substructure_data.
@@ 4.1 Protostellar disk substructures versus $T_{\rm bol}$ and disk properties
Figure 1: Protostellar disk substructure detections as a function of disk radius, mass, and bolometric temperature. The latter is taken from the eHOPS catalog. The gray dashed lines indicate the approximate boundary between the Class 0 and Class I, and between the Class I and Class II evolutionary phases (Chen et al., 1995), along with the corresponding expected duration of each Class (Dunham et al., 2015). While the $T_{\rm bol}$ is not linear with age, in general the Class 0 sources have $T_{\rm bol}<70$ K, Class I sources have $65070<T_{\rm bol}<$ 650 K, and Class II and Class III sources have $T_{\rm bol}>650$ K (Bouvier et al., 2021). The black arrow is used to indicate that the particular source has also been classified as Class II in the literature (Narayanan et al., 2023).
@@
Figure 2: Substructure detection as a function of effective resolution, 1.3 mm flux and bolometric temperature, the latter from the eHOPS catalog. The black arrow is used to indicate that the particular source has also been classified as Class II in the literature (Narayanan et al., 2023).
@@
Figure 3: The upper limit of the onset of disk substructures. Same as Figure 1, but the bolometric temperature is derived from the SED fitting without the Herschel Space Observatory data ($70-500\,\mu$m) (Dunham et al., 2015). Extinction correction is applied to the infrared data, which further increases the bolometric temperature (Dunham et al., 2015). The effective resolution and the ALMA Band 6 1.3 mm flux of protostellar disks are color-coded and shown in different symbols according to the presence or absence of substructure. Compact disks with diameters smaller than 4 beams are indicated by crosses. The black arrow is used to indicate that the particular source has also been classified as Class II in the literature (Narayanan et al., 2023).
@@
Figure 4: The upper limit of the onset of disk substructures. Same as Figure 2, but the bolometric temperature is derived from the SED fitting without the Herschel Space Observatory data ($70-500\,\mu$m) (Dunham et al., 2015). Extinction correction is applied to the infrared data, which further increases the bolometric temperature (Dunham et al., 2015). The crosses represent compact disks that are not well-resolved with diameters less than 4 times the angular resolution of the observation (beams). The lower limit of the onset of the disk substructures at 200 K found in Figure 1 is also labeled. The black arrow is used to indicate that the particular source has also been classified as Class II in the literature (Narayanan et al., 2023).
@@
Figure 1 illustrates the radius and mass of the protostellar disks as a function of bolometric temperature ($T_{\rm bol}$) derived from the eHOPS catalog. The black dashed lines indicate the $T_{\rm bol}$ that marks the approximate boundary between the Class 0 and Class I, and between the Class I and Class II evolutionary phases (Chen et al., 1995), along with the corresponding expected duration of each Class (Dunham et al., 2015). For reference, we also include the sources that do not comply with the effective resolution criteria ($\theta_{D}/\theta_{\rm res}\geq 4$). The latter sources have, in general, smaller sizes and fluxes/masses.
@@
We find no detection of substructure in any of the Class 0 disks in the sample. Figure 2 shows the effective resolution and flux of the disks with and without substructures as a function of bolometric temperature, color-coded by the linear resolution. We can observe that the lack of substructure in Class 0 disks is evident across resolutions from 4 to 18 au.
Figure 3 and 4 show similar plots as Figure 1 and 2 but considering the upper limit to bolometric temperatures from Dunham et al. (2015). In that case, the earliest substructures discussed above correspond to $T_{\rm bol}$ of 163 and 300 K, for IRAS 04169+2702666For IRAS 04169+2702, the $T_{\rm bol}$ is adopted from Ohashi et al. (2023). and CFHTWIR-Oph79, respectively.
@@ 4.2 Disk substructure detection rates
Figure 5: Disk substructure detection rates versus the evolutionary stage traced by bolometric temperature. The blue and orange represent the lower and upper limits with the bolometric temperature ($T_{\rm bol}$) measured by the eHOPS (with Herschel, but not extinction corrected) and by Dunham et al. (2015) (no Herschel, but extinction corrected), respectively. The bin size of 8 disks is used to compute the substructure detection rates. The shaded region represents the $1\sigma$ uncertainty from counting statistics. All 51 protostellar disks are observed by ALMA Band 6 at 1.3 mm wavelength, have a disk diameter greater than 4 times the resolution ($\theta_{D}/\theta_{res}\geq 4$), with an inclination angle less than 75 , and linear resolution less than 20 au.
We present the protostellar disk substructure detection rates as a function of bolometric temperature in Figure 5. Each disk is assigned a value of 1 if it exhibits substructures and 0 if it does not. We compute a moving average with a window size of 8, where the kernel slides through $T_{\rm bol}$ and calculates the mean for every group of 8 disks. This means represents the substructure detection rate. Due to the fixed window size, the detection rates are sampled at discrete levels of 0/8, 1/8, 2/8, …, up to 8/8 — effectively producing 8 bins in the substructure detection rate space.
All 51 protostellar disks in Figure 5 are observed by ALMA Band 6 at 1.3 mm wavelength, have a disk diameter greater than 4 times the resolution ($\theta_{D}/\theta_{res}\geq 4$), with an inclination angle less than 75 , and linear resolution less than 20 au. We adopt the moving average with a window size of 8 to ensure enough sampling resolution across $T_{\rm bol}$ and substructure detection rate space. The blue and orange represent the case when the bolometric temperature is adopted from eHOPS catalog and the $T_{\rm bol}$ from Dunham et al. (2015), respectively. The shaded region represents the uncertainty from the binomial counting uncertainty ($\sigma_{b}$), which is calculated as
@@
Figure 6: The continuum images of “C-type” and “F-type” disks as a sequence of evolutionary class. Both types of disk substructures can be detected throughout the Class I to Class II phases.
Figure 7: Dust continuum images of disks with multiple rings in our sample and the literature. Most of the multiple-ring systems are centrally filled (F-type) (Andrews et al., 2018a; Cieza et al., 2021; Ohashi et al., 2023; Hsieh et al., 2024; Cieza et al., 2019; ALMA Partnership et al., 2015; Stephens et al., 2023). The lack of a large central cavity ($\geq$ 25 % of the disk diameter) implies that a centrally filled disk is a general feature of multiple-ring systems. The cyan circle represents the beam size. The white lines mark the scale of 50 au. Note that HD 163296, RU Lup, AS 209, GW Lup, HD 142666, HD 143006, and Sz 129 either do not have reliable bolometric temperature ($T_{\rm bol}$) measurements or their $T_{\rm bol}\geq 1900$ K, thus are not included in this study.
Figure 8: Upper panel: Protostellar disk substructure with or without a central cavity as a function of disk radius, mass, and bolometric temperature. Middle panel: Fraction of disk with a central cavity versus the evolutionary stage traced by bolometric temperature. Bottom panel: Fraction of disk with a central cavity versus the disk mass and radius. The bin size of 8 disks is used to compute the fraction. The shaded region represents the $1\sigma$ uncertainty from counting statistics. All 24 protostellar disks are observed by ALMA Band 6 at 1.3 mm wavelength, have a disk diameter greater than 4 times the resolution ($\theta_{D}/\theta_{res}\geq 4$), with an inclination angle less than 75 , have substructures, and linear resolution less than 20 au.